OpenStudy (anonymous):

If 1 equals with 2, 4 equals with 20, and 9 equals with 90, then 10 equals with what?

Parth (parthkohli):

1 => 1 * (1 + 1) => 2 4 => 4 * (4 + 1) => 20 Hint.

OpenStudy (anonymous):

2=1(1+1) 20=4(4+1) 90=9(9+1) so for 10, 10(10+1)=answer!